# Kinematic censorship and high energy particle collisions in the Schwarzschild background

### v2

@@ Abstract
We consider near-horizon collisions between two particles moving freely in the Schwarzschild metric in the region outside the horizon. One of them emerges from a white  hole. We scrutiny when such a process can lead to the indefinitely large growth of the energy in the center of mass frame in the point of collision. We also trace how the kinematics of collision manifests itself in preserving the principle of kinematic censorship according to which the energy released in any event of collision cannot be literally infinite. According to this principle, the energy released in any event of collision, must remain finite although it can be made as large as one likes. Also, we find that particle decay near the singularity leads to unbounded release of energy independently of its initial value.
@@ I Introduction
High-energy particle collisions near black holes became a hot topic after an observation made in ban . If two particles fall towards the rotating black hole and collide near its horizon, under certain conditions the energy $E_{c.m.}$ in the center of mass frame can grow unbounded. This is called the BSW effect (after the names of its authors). It was found for extremal rotating black holes and requires one of particles to be fine-tuned, $E=\omega_{H}L$ ban , prd . Here, $E$ is the Killing energy, $L$ being the angular momentum, $\omega_{H}$ the angular speed of a black hole. For the Schwarzschild black hole this is impossible since $\omega_{H}=0$. Meanwhile, there is another version (or analogue) of the BSW effect, valid even in the Schwarzschild background. It implies head-on collision between particles, when, say, particle 1 approaches the horizon while particle 2 moves away from  it. In doing so, no fine-tuning is required, $E_{c.m.}$ grows indefinitely for any individual energies. As a particle cannot emerge from the black hole horizon, this scenario implies the presence of a white hole. Collisions of such a type were considered in gpwhite , white . Before, the fact that collision between particles with opposite radial velocity leads to unbounded $E_{c.m.}$ was noticed in pir1 - pir3 . However, it was not pointed out there that any reasonable physical realization of such a scenario corresponds just to white (not black) hole horizons. A coherent scheme of possible high energy particle collisions involving particles that emerge from a white hole, was built in frontal depending on possible fine-tuning parameters of one of two particles.
Meanwhile, in the present article, we consider other aspects of head-on collisions. We take a simplest and physically relevant metric - the Schwarzschild one. Then, all particles are not fine-tuned, so we consider collisions between two usual particles. This corresponds to line 1 in Table 1 of Ref. frontal . Thus we restrict ourselves by one type of collision only but for this type scrutinize all possible types of trajectories in space-time and their kinematic properties. In this context, we make special emphasis on manifestation of the principle of kinematic censorship cens . According to this principle, in any event of collision, only finite energy can be released. If some scenario admits literally infinite $E_{c.m.}$, it is based on some unphysical assumption or mistake and should be rejected. Sometimes, it is quite non-trivial task to trace how this principle reveals itself in concrete scenarios gen . In the present work, we take a simplest case and consider collisions in our region (outside the horizon). Even in this case, corresponding explanation, how this principle reveals itself, is quite involved and relies heavily on some subtle details of particle kinematics.
@@
The paper is organized as follows. In Sec. II we give the expression for the metric, equations of motion of free particles and the main formulas for particle collisions. In Sec. III and IV we describe in detail main scenarios of particle collisions discussed in the present paper. In Sec. V we discuss the relations between the principle of kinematic censorship and properties of trajectories in question. In Sec. VI we discuss the properties of the Schwarzschild time for a particle moving near a white whole and relate them to our scenarios. In Sec. VII we describe these scenarios using the expanding Lemaître frame. In Sec. VIII we consider processes with particles near the singularity. In Sec. IX we give summary of the results obtained.
@@ II General setup
Hereafter, we will use, along with original Schwarzschild coordinates (1), also Kruskal-Szekeres ones since this enables us to avoid coordinate singularities on the horizon. The whole space-time contains four regions (see. e.g. Fig. 31.3 of mtw ). However, for our purposes we need to consider only two of them. These are our region outside the horizon and the white hole one $T^{+}$ under the horizon fn .
@@ III.1 Properties of the trajectory of particle 1
$r^{\ast}=\int_{r_{1}}^{r}\frac{dr^{\prime}}{f(r^{\prime})}.$ (8)
@@
$t=t_{0}+\varepsilon\int_{r}^{r_{0}}\frac{dr^{\prime}}{P(r^{\prime})f(r^{\prime})}\text{,}$ (13)
@@
As $U_{c}<0$ is arbitrary according to assumption, $u_{c}$ is finite (hereafter, subscript ”c” means that the corresponding quantity refers to the point of collision). Then, according to (7) $v_{c}$ should be large negative when $r\rightarrow r_{+}$. We assume that $r_{0}$ is fixed and finite, so $r_{0}^{\ast}$ is finite as well. Meanwhile, $t_{0}$ is a parameter allowed to vary.
If $t_{0}$ is finite, $u_{0}$ is finite as well. As the integral in (14) diverges logarithmically when $r\rightarrow r_{+}$, it means that $u\rightarrow+\infty,U\rightarrow 0$ and the point of collision is close to the future horizon instead of being close to the past one, contrary to assumption. To repair this, we assume that $u_{0}$ is big and negative to compensate the integral which is big positive. This is achieved by the choice of big negative $t_{0}$, so formally
$t_{0}\rightarrow-\infty\text{.}$ (20)
@@
$V_{c}\approx\exp(\frac{2t_{0}}{r_{+}})V^{\ast}\text{,}$ (23)
@@ III.2 Properties of the trajectory of particle 2
Iw we consider collisions and assume Scenario 1, we must take into account (23) and the fact that $U_{c}=O(1)$. Then,
$r_{c}-r_{+}\backsim\exp(\frac{2t_{0}}{r_{+}})\backsim V_{c}\text{.}$ (29)
$\gamma\backsim V_{c}^{-1}\backsim\exp(-\frac{2t_{0}}{r_{+}})\backsim\frac{1}{r_{c}-r_{+}}\text{,}$ (30)
Thus we see that the same parameter $t_{0}$ (which is large negative) governs both the behavior of the Lorentz gamma factor (hence,   $E_{c.m.}$) and the proximity of the point of collision to the horizon.
@@ IV.2 Properties of the trajectory of particle 2
In terms of the Kruskal-Szekeres coordinates, for particle 2 we have in region I
@@
$v_{0}^{(2)}=\frac{2}{\kappa}\left|\ln\frac{r_{c}-r_{+}}{r_{+}}\right|+v_{2}\text{,}$ (39)
@@
$\left|U_{p}\right|\backsim\exp(-\frac{2t_{0}^{(2)}}{r_{+}})\backsim r_{c}-r_{+}$ (40)
@@
V Divergent $E_{c.m.}$ , unattainability of the speed of light and kinematic censorship
@@ V.1 Scenario 1
For any finite $t_{0}$, the energy $E_{c.m.}$ remains finite, although it can be made as big as one likes. In the limit $t_{0}\rightarrow-\infty$ it diverges. But, clearly, the moment $t_{0}$ cannot be equal to infinity literally. To some extent, this example is complimentary to that for extremal black holes. For them, $E_{c.m.}$ is infinite if collision happens precisely on the horizon and one of colliding particles is fine-tuned. But this requires infinite proper time $\tau$ to reach the horizon ted . We can take it not exactly fine-tuned, then $\tau$ is finite. But $E_{c.m.}$ is finite as well. In our example it is coordinate (not proper) time diverges and this restricts $E_{c.m.}$
@@
There is also another interesting property of this scenario. As is mentioned above, without collision particle 1 would continue its trajectory and hit the future horizon in the point very close to the bifurcation point. The proximity to it is governed by the same parameter $t_{0}$. Formally, infinite $\left|t_{0}\right|$ would mean that particle 1 passes though the bifurcation point precisely. In turn, this would mean movement along the horizon. But a massive particle cannot move along the light-like surface, its trajectory is time-like. Thus we can say that impossibility to reach infinite $E_{c.m.}$ is connected (although indirectly) with impossibility for a massive particle to attain the speed of light.
V.2 Scenario 2
@@
In this scenario, for particle 2 the quantity $t_{0}^{(2)}\rightarrow+\infty$ along with infinite $E_{c.m.}.$ Under the horizon, as is said above, this implies infinite distance along the leg of cylinder that is impossible. Also, the same arguments as above, show that infinity $t_{0}^{(2)}$ would mean that a massive particle moves along the light-like surface. The difference between scenarios 1 and 2 consists in this context in that in the first case the past horizon is relevant whereas in the second case it is the future horizon.
@@
VI Discussion: collision near white hole and behavior of the Schwarzschild time
@@ VI.2 Particle 2
Thus it is quite typical for outgoing particle 2 to have $t_{c}$ finite. What is more unusual is that $t_{c}$ remains finite in the limit $r_{c}\rightarrow r_{+}$ for ingoing particle 1. It would seem that this is contrasted with the behavior known even from textbooks according to which $t_{c}\rightarrow\infty$ in such circumstances. However, one should bear in mind two qualitative features here: (i) now the horizon is a past horizon, not a future one, (ii) collision occurs near a white (not a black) hole. Mathematically, finite $t_{c}$ appears due to compensation of a big constant $t_{0}^{(1)}$ and big (almost diverging) integral in (13) as is explained above in III.
VII How does the picture look like in the Lemaître coordinates?
@@
Near the horizon when an initial point $r\rightarrow r_{+}+0$, the metric function $f\rightarrow 0$ and both integrals compensate each other, so $\hat{t}_{c}$ is finite for any finite $t_{0}^{(2)}$. This is quite natural as this frame is designed just for description of outgoing particles. In other words, divergent $t_{c}(r)$ for $r\rightarrow r_{+}$ is combined with the finiteness of $\hat{t}(r)$, as it should be. This agrees with the behavior of $t(r)$ described in Sec. VI.
@@
As particle 1 starts at $r=r_{0}>r_{+}$, both integrals have the same sign and diverge when $r\rightarrow r_{+}.$ More aspects of relation between different frames and particles moving in the background under discussion can be found in we2 .
@@ VII.1 Scenario 1
VII.2 Scenario 2
Now, there are no special restriction imposed on particle 1, $t_{0}^{(1)}$ is finite. As a result, for the point of collision $\hat{t}\rightarrow+\infty$. For particle 2, we choose $t_{0}^{(2)}\rightarrow+\infty$ to achieve $\hat{t}_{1}=\hat{t}_{2}$. Note that is this scenario the Lemaître time of the collision diverges for divergent energy.
VIII Particle decay near singularity

### v3

@@ Abstract
We consider near-horizon collisions between two particles moving freely in the Schwarzschild metric in the region outside the horizon. One of them emerges from a white  hole. We scrutiny when such a process can lead to the indefinitely large growth of the energy in the center of mass frame in the point of collision. We also trace how the kinematics of collision manifests itself in preserving the principle of kinematic censorship. According to this principle, the energy released in any event of collision, must remain finite although it can be made as large as one likes. Also, we find that particle decay near the singularity leads to unbounded release of energy independently of its initial value.
@@ I Introduction
High-energy particle collisions near black holes became a hot topic after an observation made in ban . If two particles fall towards the rotating black hole and collide near its horizon, under certain conditions the energy $E_{c.m.}$ in the center of mass frame can grow unbounded. This is called the BSW effect (after the names of its authors). It was found for extremal rotating black holes and requires one of particles to be fine-tuned, $E=\omega_{H}L$ ban , prd . Here, $E$ is the Killing energy, $L$ being the angular momentum, $\omega_{H}$ the angular speed of a black hole. For the Schwarzschild black hole this is impossible since $\omega_{H}=0$. Meanwhile, there is another version (or analogue) of the BSW effect, valid even in the Schwarzschild background. It implies head-on collision between particles, when, say, particle 1 approaches the horizon while particle 2 moves away from  it. In doing so, no fine-tuning is required, $E_{c.m.}$ grows indefinitely for any individual energies. As a particle cannot emerge from the black hole horizon, this scenario implies the presence of a white hole. Collisions of such a type were considered in gpwhite , white . Before, the fact that collision between particles with opposite radial velocity leads to unbounded $E_{c.m.}$ was noticed in pir1 - pir3 . However, it was not pointed out there that any reasonable physical realization of such a scenario corresponds just to white (not black) hole horizons. A coherent scheme of possible high energy particle collisions involving particles that emerge from a white hole, was built in frontal depending on possible type of trajectory.
Meanwhile, in the present article, we consider other aspects of head-on collisions. We take a simplest and physically relevant metric - the Schwarzschild one. Then, all particles are not fine-tuned, so we consider collisions between two usual particles. This corresponds to line 1 in Table 1 of Ref. frontal . Thus we restrict ourselves by one type of collision only but for this type scrutinize all possible types of trajectories in space-time and their kinematic properties. In this context, we make special emphasis on manifestation of the principle of kinematic censorship cens . According to this principle, in any event of collision, only finite energy can be released. If some scenario admits literally infinite $E_{c.m.}$, it is based on some unphysical assumption or mistake and should be rejected. Sometimes, it is quite non-trivial task to trace how this principle reveals itself in concrete scenarios gen . In the present work, we take a simplest case and consider collisions in ”our” region (outside the horizon). Even in this case, corresponding explanation how this principle reveals itself, is quite involved and relies heavily on some subtle details of particle kinematics.
@@
The paper is organized as follows. In Sec. II we give the expression for the metric, equations of motion of free particles and the main formulas for particle collisions. In Sec. III and IV we describe in detail main scenarios of particle collisions discussed in the present paper. In Sec. V we mention an additional scenario that incorporates features of Scenarios 1 and 2. In Sec. VI we discuss the relations between the principle of kinematic censorship and properties of trajectories in question. In Sec. VII we discuss the properties of the Schwarzschild time for a particle moving near a white whole and relate them to our scenarios. In Sec. VIII we describe these scenarios using the expanding Lemaître frame. In Sec. IX we consider processes with particles near the singularity. In Sec. X we give summary of the results obtained.
@@ II General setup
Hereafter, we will use both the original Schwarzschild coordinates (1) and Kruskal-Szekeres ones. The whole space-time contains four regions (see. e.g. Fig. 31.3 of mtw ). However, for our purposes we need to consider only two of them. These are our region outside the horizon and the white hole one $T^{+}$ under the horizon fn .
@@ III.1 Properties of the trajectory of particle 1
$r^{\ast}=\int_{r_{1}}^{r}\frac{dr^{\prime}}{f(r^{\prime})},$ (8)
@@
$t=t_{0}^{(1)}+\varepsilon\int_{r}^{r_{0}}\frac{dr^{\prime}}{P(r^{\prime})f(r^{\prime})}\text{,}$ (13)
@@
As $U_{c}<0$ is arbitrary according to assumption, $u_{c}$ is finite (hereafter, subscript ”c” means that the corresponding quantity refers to the point of collision). Then, according to (7), $v_{c}$ should be large negative when $r\rightarrow r_{+}$. We assume that $r_{0}$ is fixed and finite, so $r_{0}^{\ast}$ is finite as well. Meanwhile, $t_{0}^{(1)}$ is a parameter allowed to vary.
If $t_{0}$ is finite, $u_{0}$ is finite as well. As the integral in (14) diverges logarithmically when $r\rightarrow r_{+}$, it means that $u\rightarrow+\infty,U\rightarrow 0$ and the point of collision is close to the future horizon instead of being close to the past one, contrary to assumption. To repair this, we assume that $u_{0}^{(1)}$ is big and negative to compensate the integral which is big positive. This is achieved by the choice of big negative $t_{0}^{(1)}$, so formally
$t_{0}^{(1)}\rightarrow-\infty\text{.}$ (20)
@@
$V_{c}\approx\exp(\frac{t_{0}^{(1)}}{r_{+}})V^{\ast}\text{,}$ (23)
@@
To conclude this subsection, let us make some remarks of general character. Usually, properties of particle motion are insensitive to the choice of the moment $t_{0}$ of time start. The more so, this seems to be irrelevant in the static metric However, now we deal with the process in which two particles are involved, so that their trajectories should agree with each other. This makes in eq. (13) not only the $r$ dependence of the integral important but also the choice of $t_{0}$. We will see below that behavior of the constant of integration is an essential ingredient of high energy collision scenarios. And, this is valid not only for scenario 1 but for scenario 2 as well.
@@ III.2 Properties of the trajectory of particle 2
If we consider collisions and assume Scenario 1, we must take into account (23) and the fact that $U_{c}=O(1)$. Then,
$r_{c}-r_{+}\backsim\exp(\frac{t_{0}^{(2)}}{r_{+}})\backsim V_{c}\text{.}$ (29)
$\gamma\backsim V_{c}^{-1}\backsim\exp(-\frac{t_{0}^{(2)}}{r_{+}})\backsim\frac{1}{r_{c}-r_{+}}\text{,}$ (30)
Thus we see that the same parameter $t_{0}^{(2)}$ (which is large negative) governs both the behavior of the Lorentz gamma factor (hence,   $E_{c.m.}$) and the proximity of the point of collision to the horizon. To distinguish between particles, we indicate explictly their number in a superscript.
@@ IV.2 Properties of the trajectory of particle 2
In terms of the Kruskal-Szekeres coordinates, for particle 2 we have in the region outside the horzon
@@
$v_{0}^{(2)}=\frac{1}{\kappa}\left|\ln\frac{r_{c}-r_{+}}{r_{+}}\right|+v_{2}\text{,}$ (39)
@@
$\left|U_{p}\right|\backsim\exp(-\frac{t_{0}^{(2)}}{r_{+}})\backsim r_{c}-r_{+}.$ (40)
@@
V Scenario 3
Scenarios 1 and 2 look complementary to each other since in the first case $t_{0}^{(1)}\rightarrow-\infty$, $t_{0}^{(2)}$ is finite and in the second case $t_{0}^{(1)}$ is finite and $t_{0}^{(2)}\rightarrow+\infty$. Meanwhile, there is a way to combine them in another scenario. This can be achieved if simultaneously $t_{0}^{(1)}\rightarrow-\infty$ and $t_{0}^{(2)}\rightarrow+\infty$. See Fig. 3. In the Carter-Penrose diagrams trajectories of both particles go closely to two branches of the past horizon.
VI Divergent $E_{c.m.}$ , unattainability of the speed of light and kinematic censorship
@@ VI.1 Scenario 1
For any finite $t_{0}^{(1)}$, the energy $E_{c.m.}$ remains finite, although it can be made as big as one likes. In the limit $t_{0}^{(1)}\rightarrow-\infty$ it diverges. But, clearly, the moment $t_{0}$ cannot be equal to infinity literally. To some extent, this example is complimentary to that for extremal black holes. For them, $E_{c.m.}$ is formally infinite if collision would happen precisely on the horizon and one of colliding particles is fine-tuned. But this requires infinite proper time $\tau$ to reach the horizon ted . We can take it not exactly fine-tuned, then $\tau$ is finite. But $E_{c.m.}$ is finite as well. In our example it is coordinate (not proper) time which diverges and this restricts $E_{c.m.}$
@@
There is also another interesting property of this scenario. As is mentioned above, without collision particle 1 would continue its trajectory and hit the future horizon in the point very close to the bifurcation point. The proximity to it is governed by the same parameter $t_{0}^{(1)}$. Formally, infinite $\left|t_{0}^{(1)}\right|$ would mean that particle 1 passes though the bifurcation point precisely. In turn, this would mean movement along the horizon. But a massive particle cannot move along the light-like surface, its trajectory is time-like. Thus we can say that impossibility to reach infinite $E_{c.m.}$ is connected (although indirectly) with impossibility for a massive particle to attain the speed of light.
We can rephrase this in somewhat different way. We arrange collision in a point with a given $t_{c}$.If we keep $t_{c}$ fixed, when sending $t_{0}^{(1)}\rightarrow-\infty$, the point of collision becomes closer and closer to the horizon, the integral in eq. (13) grows without bound in such a way that divergent terms compensate each other. Meanwhile, in the movement in the static background (see, e.g. eq. 88.9 in LL ) the conserved energy
$E=\frac{m\sqrt{f}}{\sqrt{1-V^{2}}}\text{,}$ (43)
where $V$ is velocity measured by a static observer, the speed of light $c=1$. As particle 1 falls towards the horizon, the value of $r$ decreases from $r>r_{+}$ to $r=r_{+}$ and the horizon is reachable only in the limit. In doing so, $\dot{r}<0$ when crossing the horizon, so it cannot move along $r=const=r_{+}$. In other words, before reaching the future horizon, a particle cannot move along the past horizon. Correspondingly, collision with particle 2 cannot occur exactly on the past horizon and $\ E_{c.m.}$ cannot become infinite in agreement with the principle of the kinematic censorship. The limit $V=1$ would have been achieved for a massive particle on the future (not past) horizon, if collision had not occured. Thus three circumstances cannot be realized for a massive particle: (i) infinite $E_{c.m.}$, (ii) collision on the past horizon, (iii) exact equality $V=1$.
VI.2 Scenario 2
@@
In this scenario, for particle 2 the quantity $t_{0}^{(2)}\rightarrow+\infty$ along with infinite $E_{c.m.}.$ Under the horizon, as is said above, this implies infinite distance along the leg of cylinder that is impossible, distance can be as large as one likes but it remains finite anyway. Also, the same arguments as above, show that infinite $t_{0}^{(2)}$ would mean that a massive particle moves along the light-like surface. The difference between scenarios 1 and 2 consists in this context in that in the first case the past horizon is relevant whereas in the second case it is the future horizon.
The above discussion shows relationship between changing the constants $t_{0}^{(1)}$, $t_{0}^{(2)}$and the proximity of collision to the horizon, hence an infinite growth of $E_{c.m.}$ due to velocity $V\,$ approaching 1. By itself, the fact that from the viewpoint of a stationary observer $V\rightarrow 1$ in the horizon limit, is well known. But now, we put it in the context directly connected with the principle of kinematic censorship. One should bear  in mind that the static frame in which $V\rightarrow 1$ becomes singular on the horizon where it cannot be realized by physical bodies. The question how these issues look for a physical observers measuring velocities is  considered below in Sec. VIII. In this sense, we trace the examination of the kinematic censorship from two different points.
@@
VII Discussion: collision near white hole and behavior of the Schwarzschild time
@@ VII.2 Particle 2
Thus it is quite typical for outgoing particle 2 to have $t_{c}$ finite. What is more unusual is that $t_{c}$ remains finite in the limit $r_{c}\rightarrow r_{+}$ for ingoing particle 1. It would seem that this is contrasted with the behavior known even from textbooks according to which $t_{c}\rightarrow\infty$ in such circumstances. However, one should bear in mind two qualitative features here: (i) now the horizon is a past horizon, not a future one, (ii) collision occurs near a white (not a black) hole. Mathematically, finite $t_{c}$ appears due to compensation of a big constant $t_{0}^{(1)}$ and big (almost diverging) integral in (13) as is explained above in III. And, in any case the difference in time $t_{c}-t_{0}$ diverges when a point of collision approaches the horizon.
VIII How does the picture look like in the Lemaître coordinates?
@@
Near the horizon when $r_{c}\rightarrow r_{+}+0$, the metric function $f\rightarrow 0$ and both integrals in (48) compensate each other, so $\hat{t}_{c}$ is finite for any finite $t_{0}^{(2)}$. This is quite natural as this frame is designed just for description of outgoing particles. In other words, divergent $t_{c}(r)$ for $r\rightarrow r_{+}$ is combined with the finiteness of $\hat{t}(r)$, as it should be. This agrees with the behavior of $t(r)$ described in Sec. VII.
@@
As particle 1 starts at $r=r_{0}>r_{+}$, both integrals have the same sign and diverge when $r\rightarrow r_{+}.$ More aspects of relation between different frames and particles moving in the background under discussion can be found in we2 . The definition of constants in (45) and (50) for particles 1 and 2 do not coincide. This is quite natural since their trajectories are qualitatively different. Tracing back in time, an outgoing particle 2 approaches $r_{c}$ slightly above the horizon, an ingoing particle 1 approaches $r_{0}$.
@@ VIII.1 Scenario 1
It is known that the velocity of an ingoing particle with respect to expanding Lemaître frame (as well as the velocity of an outgoing particle with respect to contracting Lemaître frame) reaches the velocity of light at the horizon we2 (especially see Table 2 there). This leads us again to the connection between divergent collision energy and a massive particle with the velocity of light. Indeed, the velocity with respect to Lemaître frame can be measured directly by free moving particle since this frame is a free moving frame. So that, an ordinary particle crossing a white hole horizon outward at some time $\hat{t}_{c}$ would be able to measure a velocity of an ingoing particle colliding with it at a horizon and get the embarassing result $v=c$! What prevents such a situation is the fact that the expanding Lemaître time for an ingoing particle to reach the horizon is infinite, so that it should start motion in infinitely remote past to compensate diverging integral in (50).
VIII.2 Scenario 2
Now, there are no special restriction imposed on particle 1, $t_{0}^{(1)}$ is finite. As a result, for the point of collision $\hat{t}\rightarrow+\infty$. For particle 2, we choose $t_{0}^{(2)}\rightarrow+\infty$ to achieve $\hat{t}_{1}=\hat{t}_{2}$. Note that in this scenario the Lemaître time of the collision diverges for divergent energy.
IX Particle decay near singularity